OpenStudy (anonymous):

Given the parent function of f(x) = x4, what change will occur when the function is changed to -f(2x)?

OpenStudy (anonymous):

help please i just need an explanation on how to do this

OpenStudy (anonymous):

The function will be the reverse, if \[f(x)=x^4\]then\[-f(2x)=-(2x)^4=-16x^4\]and the size will rise.